OpenStudy (anonymous):

Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. −three fourthsx + y = 10 3x − 4y = −40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

\[y-4=\frac{3}{4}(x+8)\]?

OpenStudy (anonymous):

so where do i start.

OpenStudy (anonymous):

multiply on the right first using the distributive law you get \[y-4=\frac{3}{4}x+6\]

OpenStudy (anonymous):

then add 4 to both sides to get \(y\) by itself on the left

OpenStudy (anonymous):

so i x's 6 with 3& 4